The owner of a store sells raisins for $3.20 per pound and nuts for $2.40 per pound. He decides to mix the raisins and sell 50lb of mixture for $2.72 per pound. What quantities of raisins and nuts should he use?
Answer by htmentor(1343)

The owner of a store sells raisins for $3.20 per pound and nuts for $2.40 per pound. He decides to mix the raisins and sell 50lb of mixture for $2.72 per pound. What quantities of raisins and nuts should he use?
================================================================
Let r = the quantity of raisins
Since the total mixture is 50 lb, the quantity of nuts = 50 - r
We can write the following equation for the total dollar amount:
3.20r + 2.40(50-r) = 2.72*50
Simplify and solve:
3.2r - 2.4r + 120 = 136
0.8r = 16
r = 20
So the quantities are:
Raisins- 20 lbs
Nuts- 50 - 20 = 30 lbs